# Steiner Traveling Salesman Problem with Quantum Annealing

### v3

@@ 2.1. Characteristics of the problem
Figure 1. Representation of the STSP network: the yellow circles represent the $V_{R}$ nodes, the green circle indicates the depot and the blue circles represent the steiner nodes. The black edges represent the arcs of the graph, while a possible route is highlighted in red, showing one of the potential solutions to the optimization problem.
[…]
Graph representation of the STSP network. The graph consists of yellow, blue, and green circles connected by black and red directed edges. The yellow circles represent the $V_{R}$ nodes, the green circle is the depot, and the blue circles are the Steiner nodes. The black edges represent the arcs of the graph, while a possible route is highlighted in red, illustrating one potential solution to the optimization problem.
@@
$A$ Set of directed arcs in the graph
@@ 3. Solution approach
$|V|$ SILP Objective Function SILP Time RILP Objective Function RILP Time
@@ 4. Computational study
Experimental tests are conducted within a 100x100 unit square region. The set of nodes $V$ is randomly generated by assigning each node random Cartesian coordinates within this area. An additional node, denoted as the depot node 00, is generated with random coordinates within the same region. The subset of required nodes $V_{R}\subseteq V$ is selected comprising the first $70\%$ of the total vertices in $V$, rounded down to the nearest integer. Furthermore, each arc cost $c_{k}$ is assigned a random value between 20 and 50. To facilitate the replication of this study, all instances used are publicly accessible in (Ciacco et al., 2025b).
@@
$|V|$ QA with QPU for SQUBO Objective Function QA with QPU for SQUBO Time QA with QPU for RQUBO Objective Function QA with QPU for RQUBO Time
@@
$|V|$ QA with LeapBQM for SQUBO Objective Function QA with LeapBQM for SQUBO Time QA with LeapBQM for RQUBO Objective Function QA with LeapBQM for RQUBO Time

### v4

@@ 2.1. Characteristics of the problem
Figure 1. Representation of the STSP network: the yellow circles represent the $V_{R}$ nodes, the green circle indicates the depot and the blue circles represent the steiner nodes. The black edges represent the arcs of the graph, while a possible route is highlighted in red, showing one of the potential solutions to the optimization problem.:
@@
AA Set of directed arcs in the graph
@@ 3. Solution approach
$|V|$
SILP
Objective Function
SILP
Time
RILP
Objective Function
RILP
Time
@@ 4. Computational study
Experimental tests are conducted within a 100x100 unit square region. The set of nodes $V$ is randomly generated by assigning each node random Cartesian coordinates within this area. An additional node, denoted as the depot node 0, is generated with random coordinates within the same region. The subset of required nodes $V_{R}\subseteq V$ is selected comprising the first $70\%$ of the total vertices in $V$, rounded down to the nearest integer. Furthermore, each arc cost $c_{k}$ is assigned a random value between 20 and 50. To facilitate the replication of this study, all instances used are publicly accessible in (Ciacco et al., 2025b).
@@
$|V|$
QA with QPU for SQUBO
Objective Function
QA with QPU for SQUBO
Time
QA with QPU for RQUBO
Objective Function
QA with QPU for RQUBO
Time
@@
$|V|$
QA with LeapBQM for SQUBO
Objective Function
QA with LeapBQM for SQUBO
Time
QA with LeapBQM for RQUBO
Objective Function
QA with LeapBQM for RQUBO
Time